OpenStudy (anonymous):

simplify r^2-4/(r+2)^2

OpenStudy (anonymous):

Use the identity : a^2-b^2=(a+b)(a-b) On the top part of the fraction

OpenStudy (anonymous):

assuming the question is: (r^2-4)/((r+2)^2)

OpenStudy (lalaly):

\[\frac{r^2-4}{(r+2)^2} = \frac{(r+2)(r-2)}{(r+2)^2}\]now cancel out the r+2 \[= \frac{r-2}{r+2}\]